# Secure authentication via Quantum Physical Unclonable Functions: a review

### v2

@@ 1 Introduction
This article aims to provide the reader with a review on QPUFs, going in detail over all the aforementioned aspects within the complex development and evolution of an itself complex field of investigation. In Section 2.1, we revisit the theoretical framework introduced in [18], and we comment on some immediate implications of it when accepting certain assumptions commonly made for PUFs. In Section 2.2, we deliver a discussion on the current incompatibility between requirements and experimental possibilities for QPUF-based tokens, as well as the corresponding near-term perspectives. Later, in Section 3, we present and justify the article selection criteria used in this review, focusing on the most theoretically relevant QPUF proposals. In Section 4, we present our analysis of several articles, providing an underlying historical overview of QPUFs—illustrated as a QPUF timeline—, and a detailed discussion on information-theoretic approaches used for the study and characterization of QPUFs. Section 5 provides an outlook on future research and Section 6 concludes the article.
@@ 2. Response as a token Holds the QPUF Stores a response (token) 1-shot verification
As already pointed out in [18], we notice how the assumption made on QPUF unclonability could be violated via process tomography [32] within the, also assumed, query-based adversary model. In this context, exponential unforgeability is unattainable. However, an appropriate choice of security parameters, e.g. the number of qubits targeted by the QPUF channel and the number of responses requested during verification, can potentially lead to achieving the other two notions of unforgeability, which are more realistic and typically sufficient. Importantly, as we discuss in detail in Section 4, we thus notice how Definition 2.1 denies the quality of being a QPUF for the QR-PUF models [19, 20, 33, 34, 35], which target systems of fixed size thus leading to learnable quantum channels unless further assumptions are considered.
@@
As stated in Remark 2.2, a quantum-channel sampling procedure must be available at the verifier side and, informally, it should not allow for having two too similar instances among different runs of the QPUF-generation scheme. Now, additionally, the unclonability assumption implicitly requires such sampling procedure to be non-reproducible. The security proof (Theorem 6 in [18]) for the weakest form of QPUF unforgeability, i.e., the selective one, motivates the cited authors to model such sampling procedure as being Haar random. That is, Haar randomness paves the way towards having robust security guarantees. Nonetheless, to the best of our knowledge, there is currently no efficient method for mimicking such mathematical construction in real setups. Specifically, one can see how the two prescriptions for Haar-randomly sampling unitary operators given in [24] introduce an exponential overhead in the number of targeted qubits.
@@ 3 Methodology
In this review, we have selected the works of QR-QPUF [101], QPUF [18, 26], and Hybrid PUFs (HPUFs) [17] as our baseline articles, as they introduce novel theoretical perspectives related to the QPUF topic. It is important to note a selection bias in our choice of baselines, as we primarily focus on approaches that emphasize theoretical protocol innovation. Furthermore, for our analysis, we have set the acceptance threshold value to $l=0.1$. The outcome of the article selection is presented in Table  5 within the Appendix, where baseline articles are highlighted in yellow, and accepted articles for review are shaded in light gray. To better illustrate the relationships between articles, Figure  1 shows the co-citation network, with the article [18] serving as a baseline reference.
@@ 4.2 QPUFs
    Secondly, the work in [31] makes different relevant contributions to the field. On the one hand, the cited authors note that [18] lacks a uniqueness proof for their QPUF theoretical construction. Nevertheless, they show that uniqueness is guaranteed by the Haar-randomness hypothesis. Notably, they additionally prove that other sampling strategies can also deliver uniqueness. Furthermore this article comments on the inconvenience of the Haar-randomness requirement, and proposes an alternative relying on the well-known properties of $t$-designs, delivering the first application of this concept for a general value of $t$. The proper functioning of this alternative, however, comes at the price of restricting the number of queries by the adversary to $t$, instead of it being any polynomial amount of certain security parameters. As a final remark, this work does not omit a discussion on the effect of quantum noise, but they restrict it to the case of unitary noise channels. In such case the proposed scheme remains functional.
@@ 6 Conclusions
Before concluding this article, we want to further notice how, in this work, we have discussed how some quantum-based PUFs—such as QR-PUFs—can still be regarded as partly classical. Interestingly, the converse also holds: certain commercial implementations of classical PUFs, such as those by Crypto Quantique [124] and PUFsecurity [125], rely on fabrication-level variations originating from quantum phenomena, including tunneling in Complementary Metal–Oxide–Semiconductor (CMOS) or Metal–Oxide–Semiconductor Field-Effect Transistor (MOSFET) devices. Moreover, across all forms of PUFs, ensuring Truly Random Generation appears essential for security and unpredictability, as highlighted in recent industrial developments integrating quantum-based randomness [126, 127]. Quantum technologies can contribute significantly to this critical cryptographic primitive [128, 129].

### v3

@@ 1 Introduction
This article aims to provide the reader with a review on QPUFs, going in detail over all the aforementioned aspects within the complex development and evolution of an itself complex field of investigation. In Section 2.1, we revisit the theoretical framework introduced in [18], and we comment on some immediate implications of it when accepting certain assumptions commonly made for PUFs. In Section 2.2, we deliver a discussion on the current incompatibility between requirements and experimental possibilities for QPUF-based tokens, as well as the corresponding near-term perspectives. Later, in Section 3, we present and justify the article selection criteria used in this review, focusing on the most theoretically relevant QPUF proposals. In Section 4, we present our analysis of several articles, providing an underlying historical overview of QPUFs—illustrated as a QPUF timeline—and a detailed discussion on information-theoretic approaches used for the study and characterization of QPUFs. Section 5 provides an outlook on future research and Section 6 concludes the article.
@@ 2. Response as a token Holds the QPUF Stores a response (token) 1-shot verification
As already pointed out in [18], we notice how the assumption made on QPUF unclonability could be violated via process tomography [32] within the, also assumed, query-based adversary model. In this context, exponential unforgeability is unattainable. However, an appropriate choice of security parameters, e.g. the number of qubits targeted by the QPUF channel and the number of responses requested during verification, can potentially lead to achieving the other two notions of unforgeability, which are more realistic and typically sufficient. Importantly, as we discuss in detail in Section 4, we thus notice how Definition 1 denies the quality of being a QPUF for the QR-PUF models [19, 20, 33, 34, 35], which target systems of fixed size thus leading to learnable quantum channels unless further assumptions are considered.
@@
As stated in Remark 2.2, a quantum-channel generation procedure must be available at the verifier side and, informally, it should not allow for having two too similar instances among different runs of the QPUF-generation scheme. Now, additionally, the unclonability assumption implicitly requires such generation procedure to be non-reproducible. The security proof (Theorem 6 in [18]) for the weakest form of QPUF unforgeability, i.e., the selective one, motivates the cited authors to model such generation procedure as being Haar random. That is, Haar randomness paves the way toward having robust security guarantees. Nonetheless, to the best of our knowledge, there is currently no efficient method for mimicking such mathematical construction in real setups. Specifically, one can see how the two prescriptions for Haar- randomly generating unitary operators given in [24] introduce an exponential overhead in the number of targeted qubits.
@@ 3 Methodology
In this review, we have selected the works of QR-QPUF [101], QPUF [18, 26], and Hybrid PUFs (HPUFs) [17] as our baseline articles, as they introduce novel theoretical perspectives related to the QPUF topic. It is important to note a selection bias in our choice of baselines, as we primarily focus on approaches that emphasize theoretical protocol innovation. Furthermore, for our analysis, we have set the acceptance threshold value to $l=0.1$. The outcome of the article selection is presented in Table  5 within the Appendix, where baseline articles are highlighted in yellow, and accepted articles for review are shaded in light gray. To better illustrate the relationships between articles, Figure  1 shows a co-citation network, with the article [18] serving as a baseline reference.
@@ 4.2 QPUFs
    Secondly, the work in [31] makes different relevant contributions to the field. On the one hand, the cited authors note that [18] lacks a uniqueness proof for their QPUF theoretical construction. Nevertheless, they show that uniqueness is guaranteed by the Haar-randomness hypothesis. Notably, they additionally prove that other generation strategies can also deliver uniqueness. Furthermore this article comments on the inconvenience of the Haar-randomness requirement, and proposes an alternative relying on the well-known properties of $t$-designs, delivering the first application of this concept for a general value of $t$. The proper functioning of this alternative, however, comes at the price of restricting the number of queries by the adversary to $t$, instead of it being any polynomial amount of certain security parameters. As a final remark, this work does not omit a discussion on the effect of quantum noise, but they restrict it to the case of unitary noise channels. In such case the proposed scheme remains functional.
@@ 6 Conclusions
Before concluding this article, we want to further notice how, in this work, we have discussed how some quantum-based PUFs—such as QR-PUFs—can still be regarded as partly classical. Interestingly, the converse also holds: certain commercial implementations of classical PUFs, such as those by Crypto Quantique [124] and PUFsecurity [125], rely on fabrication-level variations originating from quantum phenomena, including tunneling in Complementary Metal–Oxide–Semiconductor (CMOS) or Metal–Oxide–Semiconductor Field-Effect Transistor (MOSFET) devices. Moreover, across all forms of PUFs, ensuring True Random-Number generation appears essential for security and unpredictability, as highlighted in recent industrial developments integrating quantum-based randomness [126, 127]. Quantum technologies can contribute significantly to this critical cryptographic primitive [128, 129].